OpenStudy (anonymous):

prove that Sn = 1/1! + 1/2! + ... + 1/n! is a contractive sequence prove that Sn = 1/1! + 1/2! + ... + 1/n! is a contractive sequence @Mathematics

OpenStudy (anonymous):

so man this is over my head

OpenStudy (anonymous):

sorry*

OpenStudy (anonymous):

np, thank for looking

OpenStudy (anonymous):

Sn = 1/1! + 1/2! + ... + 1/n! Sn = 1 + 1/2 + 1/6 + 1/24 + 1/120 + 1/720 + ..............+1/n! as u can see that the denominator is increasing so the value of the first term is the highest it is decresing along the sequence so this sequence is a contractive one !

OpenStudy (anonymous):

it's not a proof, proving as |Sn+2 - Sn+1| <= K|Sn+1 - Sn|

OpenStudy (anonymous):

hahahaha................

OpenStudy (anonymous):

0 < K < 1

OpenStudy (anonymous):

Sn+1-Sn = 1/(n+1)! Sn+2 - Sn+1 = 1/(n+2)! now just plug in n= 1in the above two equations u will get Sn+1-Sn = 1/(1+1)! = 1/2! = 1/2 = 0.5 Sn+2 - Sn+1 = 1/(1+2)! = 1/3! = 1/6 = 0.16666 so\[\S_{n+2} - \S_{n+1} \le \S_{n+1} - \S_{n}\]

OpenStudy (anonymous):

it's still not a proof, it;s algebra

OpenStudy (anonymous):

oh ok. ya I will do that

OpenStudy (anonymous):

thanks